# M2IST: Multi-Modal Interactive Side-Tuning for Efficient Referring Expression Comprehension

### v2

@@ III-C Discussion: Advantages of M2IST
TABLE I: Comparison with PETL methods using the same base architecture. "Param." indicates the number of tunable parameters in the pre-trained encoders. "Mem." denotes the peak GPU memory footprint with batch size 64 during fine-tuning.
We highlight the best and the second results.
(M) (GB) val testA testB val testA testB val-g val-u test-u
@@ IV-A Datasets and Evaluation Metrics
Following previous work [3, 25, 26], we conduct experiments on both RefCOCOg-g (val-g) and RefCOCOg-u (val-u and test-u). We use Precision@0.5 as the evaluation metric. In addition to accuracy, we also report the number of tunable parameters in the pre-trained encoders and the training memory consumption in Gigabytes (GB) to compare the fine-tuning efficiency with other PETL methods.
@@ IV-C Main Results
Comparison with PETL Methods
We first compare our proposed M2IST method with several existing parameter-efficient transfer learning (PETL) approaches, using the same base architecture and the same bottleneck dimensions for the adapter modules ($C_{d}=128$). These PETL methods include vanilla Adapter [12], LoRA [13], AdaptFormer [15], CM Adapter [40], and MRS-Adapter [11].
From Table I, we can see that all PETL methods including M2IST achieve significant parameter efficiency during fine-tuning by reducing the number of tunable parameters compared to full fine-tuning. Despite the parameter efficiency achieved by these PETL methods, they still face two major limitations in practical applications: a) All methods exhibit performance degradation compared to full fine-tuning. b) They require substantial GPU memory during fine-tuning, thus failing to deliver the efficiency that is crucial during implementation.
As for the performance, M2IST stands out as the only PETL method that can achieve performance on par with or even better than full fine-tuning, significantly outperforming other PETL methods across all three benchmarks. This highlights the effectiveness of M3ISAs in adapting pre-trained knowledge for the REC task. We can observe that all PETL methods exhibit distinct performance degradation compared to full fine-tuning on the RefCOCOg dataset. This is because RefCOCOg is a substantial dataset with sufficient data, reducing the likelihood of overfitting when fully fine-tuning the models. Even so, with the facilitation of cross-modality interaction between the encoders, M3ISAs are able to enhance the modeling of complex spatial relationships, leading to competitive performance with full fine-tuning on the RefCOCOg benchmark.
Regarding fine-tuning efficiency, M2IST requires the least training memory among PETL methods. This results from the fact that gradients backpropagate through the lightweight M3ISAs rather than the heavy encoders, highlighting M2IST’s advantage in memory efficiency, as mentioned in Section III-C.
@@
Speed Full Adapter M2IST
Training (min/epoch) $\downarrow$ 52 38 33
Inference (s) $\downarrow$ 145 146 145
@@
We present the training and inference speeds on RefCOCO dataset of full fine-tuning, standard adapter-tuning [12], and our M2IST in Table III.
@@
IV-D Ablation Study and Analysis
Effects of Different Components of M3ISA
@@ IV-D Ablation Study and Analysis
In Table VI, to further investigate the effects of different adapter combination forms (i.e., mixing strategies), we present the performance of adopting intra-modality adapters or inter-modality adapters in parallel with different pre-trained layers (MHA and FFN). The findings are as follows: (1) Transferring pre-trained single-modality knowledge to the REC domain (e.g., LEA+VEA) is more effective in accurately locating the referred object than merely achieving cross-modality interaction (e.g., IEA+IEA) (Table VI (a,b)). (2) Combining intra-modality adapters and inter-modality adapters enhances performance, indicating that joint transfer of pre-trained single-modality knowledge and cross-modality interaction aids in accurately localizing referred objects by text descriptions (Table VI (a,b,c,d)). This observation aligns with findings from other challenging vision-language tasks [16, 8], suggesting that combining deep inter-modality fusion with intra-modality adaptation improves performance. (3) The best performance among the M3ISA variants is achieved by first connecting the vision and language encoders with IEAs, and then adapting the interacted features and single-modality features to the REC domain with VEA and LEA (Table VI (a,b,c,d)).
@@
As depicted in Figure 3 and Table VII, we evaluate the impact of integrating M3ISAs with different insertion forms on performance and GPU memory usage. From Table VII, we can observe that: (1) Side insertion yields the best performance. We suppose that implementing M3ISAs on side networks enhances the alignment between the referring sentence and the referred object, resulting in improved localization performance. (2) In terms of fine-tuning efficiency, all three insertion forms contribute to a reduction in GPU memory usage to varying degrees. It is evident that incorporating M3ISAs into the side networks consumes the least amount of GPU memory. This is because the gradients backpropagate through the lightweight M3ISAs instead of heavy encoders. This aligns with the memory efficiency advantage mentioned in Section III-C.
@@
As illustrated in Table VIII, we further investigate the impact of introducing M3ISAs at different positions within the pre-trained Vision Encoder and Language Encoder. The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively, and the IEA needs to be inserted into the encoder layers at the same indices. We explore three common insertion forms, as shown in Table VIII (a-c). It is evident that inserting M3ISAs in parallel to the deeper encoder layers of the pre-trained Language Encoder results in better performance. We suggest that deeper encoder layers contain richer semantic features, and establishing cross-modality interaction on this basis helps the model learn finer region-text alignment, thereby achieving better localization performance.
TABLE IX: Effects of different interaction dimensions of M3ISA. "$C_{i}$" represents the interaction dimension of Interaction Expert Adapter (IEA) in M3ISAs.
@@
Effects of Different Interaction Dimensions of M3ISA
We further ablate the impact of changing the interaction dimensions $C_{i}$ of inter-modality adapters (i.e., IEA), and follow the paradigm of Table VI (d). As depicted in Table IX, deeper cross-modality interaction provides better vision-language channel-level alignment, thus resulting in an increase in performance. Thus, $C_{i}$ is set to 256 to achieve the optimal trade-off among accuracy, number of tunable parameters, and GPU memory consumption. It is worth noting that all ablative variants exhibit a remarkable level of memory efficiency, as they consume less than 16GB of GPU memory. This observation is consistent with the memory efficiency advantage highlighted in Section III-C.
TABLE X: Effects of combining M2IST with LoRA. "Full" represents fully fine-tuning the baseline modal.
@@
As shown in Table X, using LoRA alone leads to performance degradation. This is because LoRA integrates two pairs of tunable low-rank decomposed weight matrices into each encoder layer of the Vision Encoder and Language Encoder separately, without any multi-modality interaction during the feature extraction stage. Combining M2IST with LoRA performs comparably to or even better than using M2IST alone, demonstrating the extensibility of our approach. However, as discussed in Section III-C, such a combination may significantly increase the number of updated parameters and GPU memory usage, thus undermining the memory efficiency advantage of M2IST. Therefore, our M2IST achieves the optimal performance and efficiency trade-off in this comparison.
@@ IV-E Qualitative Results
To investigate the impact of cross-modality interaction facilitated by M3ISA, we visualize the attention maps from the V-L Encoder. We compare M3ISA (referred to as "With Interaction") with its variant presented in Table VI (referred to as "Without Interaction") across various scenarios to assess their effectiveness in understanding challenging cases, such as human action recognition and spatial relation reasoning, as shown in Figure 4.
It is clear that M3ISA effectively addresses diverse REC cases, particularly in spatial relations. For instance, in the sixth case, the "With Interaction" approach successfully directs attention to the front region associated with the referring sentence "front pizza". This is especially evident in the last case, where "With Interaction" focuses more on the area corresponding to the referring sentence "right person in air", while "Without Interaction" fails to identify the correct region. These visualization results demonstrate that the enhanced cross-modality interaction facilitated by the IEA enables effective comprehension of complex semantic information.

### v3

@@ IV-A Datasets and Evaluation Metrics
Following previous work [3, 9], we conduct experiments on both RefCOCOg-g (val-g) and RefCOCOg-u (val-u and test-u). We use Precision@0.5 as the evaluation metric. In addition to accuracy, we also report the number of tunable parameters in the pre-trained encoders and the peak training memory consumption in Gigabytes (GB) to compare the fine-tuning efficiency with other PETL methods. Specifically, we consider the number of parameters in full fine-tuning as the baseline (100%) and evaluate the parameter efficiency of other methods as the percentage of their tunable parameters relative to full fine-tuning. Similarly, we consider the memory utilization of full fine-tuning as the baseline (100%) and quantify the memory efficiency of other efficient tuning methods as the ratio between their peak memory usage and that of full fine-tuning during the training phase.
@@ IV-C Main Results
Comparison with Efficient Tuning Methods
We first compare our proposed M2IST method with several existing parameter-efficient transfer learning (PETL) methods, including vanilla Adapter [20], LoRA [21], AdaptFormer [23], CM Adapter [19], and MRS-Adapter [18], as well as one memory-efficient transfer learning (METL) method, DTL [28]. All experiments use the same base architecture and maintain consistent bottleneck dimensions ($C_{d}=128$).
From Table I, we can see that all efficient tuning methods including M2IST achieve significant parameter efficiency during fine-tuning by reducing the number of tunable parameters compared to full fine-tuning. Despite the parameter efficiency achieved by these PETL methods, they still face two major limitations in practical applications: a) All methods exhibit performance degradation compared to full fine-tuning. b) They require substantial GPU memory during fine-tuning, thus failing to deliver the efficiency that is crucial during implementation.
As for the performance, M2IST stands out as the only efficient tuning method that can achieve performance on par with or even better than full fine-tuning, significantly outperforming other methods across all three benchmarks. This highlights the effectiveness of M3ISAs in adapting pre-trained knowledge for the REC task. We can observe that all PETL and METL methods exhibit distinct performance degradation compared to full fine-tuning on the RefCOCOg dataset. This is because RefCOCOg is a substantial dataset with sufficient data, reducing the likelihood of overfitting when fully fine-tuning the models. Even so, with the facilitation of cross-modality interaction between the encoders, M3ISAs are able to enhance the modeling of complex spatial relationships, leading to competitive performance with full fine-tuning on the RefCOCOg benchmark.
Regarding fine-tuning efficiency, METL methods DTL and our M2IST demonstrate superior memory efficiency compared to other PETL methods. This results from the fact that gradients backpropagate through the lightweight networks rather than the computationally intensive encoders, significantly reducing GPU memory consumption, thus highlighting M2IST’s advantage in memory efficiency, as mentioned in Section III-C. Besides, our M2IST outperforms DTL due to the fact that its core component M3ISA is able to effectively bridge pre-trained vision-language knowledge, enabling better handling of complex vision-language understanding scenarios for REC.
Speed Full Adapter M2IST
Training (min/epoch) $\downarrow$ 52 (100%) 38 (70.08%) 33 (63.46%)
Inference (s) $\downarrow$ 145 (100%) 146 (100.69%) 145 (100%)
@@
We present the training and inference speeds on RefCOCO dataset of full fine-tuning, standard adapter-tuning [20], and our M2IST in Table III, where all methods are evaluated on a single A800 GPU. For training speed, following common practice in transformer-based methods [3, 9], we train for 90 epochs on RefCOCO with a batch size of 64, and measure the average training time per epoch (min/epoch). For inference speed, we compare the inference time (s) on the RefCOCO validation set.
@@
IV-D Ablation Study and Analysis
Effects of Different Components of M3ISA
@@ IV-D Ablation Study and Analysis
In Table VI, to further investigate the effects of different adapter combination forms (i.e., mixing strategies), we present the performance of adopting intra-modality adapters or inter-modality adapters in parallel with different pre-trained layers (MHA and FFN).
The findings are as follows: (1) Transferring pre-trained single-modality knowledge to the REC domain (e.g., LEA+VEA) is more effective in accurately locating the referred object than merely achieving cross-modality interaction (e.g., IEA+IEA) (Table VI (a,b)). (2) Combining intra-modality adapters and inter-modality adapters enhances performance, indicating that joint transfer of pre-trained single-modality knowledge and cross-modality interaction aids in accurately localizing referred objects by text descriptions (Table VI (a,b,c,d)). This observation aligns with findings from other challenging vision-language tasks [24, 8], suggesting that combining deep inter-modality fusion with intra-modality adaptation improves performance. (3) The best performance among the M3ISA variants is achieved by first connecting the vision and language encoders with IEAs, and then adapting the interacted features and single-modality features to the REC domain with VEA and LEA (Table VI (a,b,c,d)).
@@
As depicted in Fig. 3 and Table VII, we evaluate the impact of integrating M3ISAs with different insertion forms on performance and GPU memory usage.
From Table VII, we can observe that: (1) Side insertion yields the best performance. We suppose that implementing M3ISAs on side networks enhances the alignment between the referring sentence and the referred object, resulting in improved localization performance. (2) In terms of fine-tuning efficiency, all three insertion forms contribute to a reduction in GPU memory usage to varying degrees. It is evident that incorporating M3ISAs into the side networks consumes the least amount of GPU memory. This is because the gradients backpropagate through the lightweight M3ISAs instead of heavy encoders. This aligns with the memory efficiency advantage mentioned in Section III-C.
@@
As illustrated in Table VIII, we further investigate the impact of introducing M3ISAs at different positions within the pre-trained Vision Encoder and Language Encoder.
The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively, and the IEA needs to be inserted into the encoder layers at the same indices. We explore three common insertion forms, as shown in Table VIII (a-c). It is evident that inserting our M3ISAs in parallel to the deeper encoder layers of the pre-trained Language Encoder results in better performance. We suggest that deeper encoder layers contain richer semantic features, and establishing cross-modality interaction on this basis helps the model learn finer region-text alignment, thereby achieving better localization performance. Therefore, we adopt insertion form (c) in practical application to achieve the optimal performance.
TABLE IX: Effects of Different Insertion Density of M3ISA. "Num." indicates M3ISA insertion count, where "0" means only updating V-L Encoder, "2" and "4" denote insertions at layers {1,4} and {1,3,4,6}, and "6" is the default setting.
# Num. Params.$\downarrow$ Mem.$\downarrow$ RefCOCO
(M) (GB) val testA testB
(a) 0 0 14.32 72.72 73.33 71.27
(b) 2 1.06 14.50 80.24 81.41 75.96
(c) 4 2.13 14.92 80.77 81.72 77.32
(d) 6 3.19 15.44 81.35 82.29 77.98
Effects of Different Insertion Density of M3ISA
We further explore the effects of M3ISA insertion density by placing M3ISAs in selected transformer encoder layers.
As presented in Table IX, from (a) and (b), we observe that incorporating M3ISA significantly improves model fine-tuning performance demonstrating effectiveness of M3ISA in adapting pre-trained vision and language models for REC tasks. Moreover, from (b,c,d), It is evident that denser integration of M3ISA modules into pre-trained networks consistently improves model performance without significantly increasing GPU memory consumption during fine-tuning, demonstrating the memory efficiency of our M2IST (see Section III-C). We adopt the default configuration of six M3ISA insertions to achieve optimal performance.
TABLE X: Effects of different interaction dimensions of M3ISA. "$C_{i}$" represents the interaction dimension of Interaction Expert Adapter (IEA) in M3ISAs.
@@
TABLE XI: Effects of combining M2IST with LoRA. "Full" represents fully fine-tuning the baseline modal. "M2IST+LoRA" represents combining M2IST with LoRA.
@@
Effects of Different Interaction Dimensions of M3ISA
We further ablate the impact of changing the interaction dimensions $C_{i}$ of inter-modality adapters (i.e., IEA), and follow the paradigm of Table VI (d).
As depicted in Table X, deeper channel-level vision-language alignment in IEA provides better cross-modality interaction, thus resulting in an increase in performance. When $C_{i}$ is set to 256 to achieve the optimal trade-off among accuracy, number of tunable parameters, and GPU memory consumption. It is worth noting that all ablative variants exhibit a remarkable level of memory efficiency, as they consume less than 16GB of GPU memory. This observation is consistent with the memory efficiency advantage highlighted in Section III-C.
@@
As shown in Table XI, using LoRA alone leads to performance degradation. This is because LoRA integrates two pairs of tunable low-rank decomposed weight matrices into each encoder layer of the Vision Encoder and Language Encoder separately, without any multi-modality interaction during the feature extraction stage. Combining M2IST with LoRA performs comparably to or even better than using M2IST alone, demonstrating the extensibility of our approach. However, as discussed in Section III-C, such a combination may significantly increase the number of updated parameters and GPU memory usage, thus undermining the memory efficiency advantage of M2IST. Therefore, our M2IST achieves the optimal performance and efficiency trade-off in this comparison.
@@ IV-E Qualitative Results
To investigate the impact of cross-modality interaction facilitated by M3ISA, we visualize the attention scores between the [REG] token and other visual tokens in the V-L Encoder, where highlighted regions represent areas with higher attention scores. We compare M3ISA (referred to as "With Interaction") with its variant presented in Table VI (referred to as "Without Interaction") across various scenarios to assess their effectiveness in understanding challenging cases, such as human action recognition and spatial relation reasoning, as shown in Fig. 4.
It is clear that M3ISA effectively addresses most of the diverse REC cases, particularly in spatial relations. For instance, in the fifth case, the "With Interaction" approach successfully directs attention to the front region associated with the referring sentence "front pizza". This is especially evident in the last case, where "With Interaction" focuses more on the area corresponding to the referring sentence "right person in air", while "Without Interaction" fails to identify the correct region. These visualization results demonstrate that the enhanced cross-modality interaction facilitated by the IEA enables effective comprehension of complex semantic information.
However, both approaches show sub-optimal performance when handling images with occluded or crowded objects, as seen in cases 3, 4, and the last example. Notably, in case 3, where the target region ("girl falling down") is occluded by another person, both variants incorrectly allocate attention to the occluding area. Similarly, in the final case ("guy in middle"), our method partially attends to neighboring regions. These cases collectively indicate limitations in local visual feature modeling."
@@ V Conclusion
VI Limitations and Future Works
In this work, we implement our M2IST on the mainstream transformer-based architecture for referring expression comprehension, comprising a pre-trained Vision Encoder and Language Encoder. However, our method shows sub-optimal performance in occluded or crowded scenes for REC. Future work will explore incorporating convolutional adapters to enhance local visual feature modeling to address such limitation.
Additionally, traditional transformer-based models demonstrate notable limitations when dealing with REC tasks in complex scenarios that demand sophisticated reasoning capabilities. Given the remarkable advancement of multi-modal large language models (MLLMs) in recent years, applying M2IST to MLLMs to models like LISA [61] and GLaMM [62] could potentially bridge this gap and significantly enhance their reasoning capabilities in complex referring scenarios. This integration would combine the efficiency of M2IST with the powerful reasoning abilities of MLLMs, creating a more robust solution for challenging REC tasks.
Furthermore, since video processing often requires more computational resources, extending our M2IST to challenging multi-modal video understanding tasks, such as referring video object segmentation [63] and video grounding [64], will be our future research direction.